# Krylov complexity as an order parameter for quantum chaotic-integrable transitions

### v1

@@ Abstract
Modern explorations of quantum chaos have prominently featured both level statistics and OTOCs, bolstered by intriguing links between many-body chaos and quantum gravitational systems Shenker and Stanford (2014, 2015); Cotler et al. (2017); de Boer et al. (2018); Stanford and Witten (2020). In this context, Krylov complexity Parker et al. (2019); Balasubramanian et al. (2022) has recently emerged as a new valuable tool for characterizing quantum chaos, providing an alternative diagnostic beyond traditional analyses. It has been employed in RMT Balasubramanian et al. (2022, 2023); Tang (2023); Caputa et al. (2024); Bhattacharjee and Nandy (2024) and many other quantum chaotic systems, including quantum billiards Hashimoto et al. (2023); Camargo et al. (2024a); Balasubramanian et al. (2024), spin chains Rabinovici et al. (2022a); Scialchi et al. (2024a); Gill et al. (2024); Bhattacharya et al. (2024); Camargo et al. (2024b); Scialchi et al. (2024b), and various flavors of the Sachdev–Ye–Kitaev (SYK) model Rabinovici et al. (2021); Bhattacharjee et al. (2023a); Hörnedal et al. (2022); Erdmenger et al. (2023a); Chapman et al. (2024). Additionally, Krylov complexity has been discussed in several other contexts including topological and quantum phase transitions Caputa and Liu (2022); Afrasiar et al. (2022); Caputa et al. (2023); Pal et al. (2023), quantum batteries Kim et al. (2022), saddle-dominated scrambling Bhattacharjee et al. (2022); Huh et al. (2024), and open quantum systems Bhattacharya et al. (2022); Bhattacharjee et al. (2023b); Mohan (2023); Bhattacharya et al. (2023); Bhattacharjee et al. (2024); Carolan et al. (2024) among others – see Nandy et al. (2024) for a comprehensive review. Two forms of Krylov complexity have been proposed in the literature: the original type, which addresses operator growth Parker et al. (2019), and a newer version, which evaluates the spread of a time-evolving quantum state within a specific subspace of the Hilbert space Balasubramanian et al. (2022). The latter will be the focus of this Letter.
@@
To provide evidence for our proposal, we turn to the SYK model and its variants, which serve as useful toy models. Specifically, we focus on the mass-deformed SYK model García-García et al. (2018), which has been extensively studied for its relevance to quantum chaotic to integrable transitions, and has been thoroughly characterized through both level statistics and properties of the OTOCs García-García et al. (2018); Nosaka et al. (2018); Kim and Cao (2021a); García-García et al. (2021); Lunkin et al. (2020); Nandy et al. (2022); Menzler and Jha (2024).
@@
Given the definitions provided in Eq.(4) and Eq.(6), it is natural to assume that a direct link between Krylov complexity and the SFF might exist. In fact, considering the Krylov complexity of the TFD state, the SFF can be interpreted as the survival probability of the time-evolved TFD state del Campo et al. (2017); Caputa et al. (2024): $\textrm{SFF}(t)=|\psi_{0}(t)|^{2}$. Additionally, for the TFD state with $\beta=0$, which is a maximally-entangled state, the late-time behavior of the Krylov complexity obeys the following identity Cotler et al. (2017); Rabinovici et al. (2021, 2022b); Erdmenger et al. (2023a)
@@
Results. We calculate the time-dependent Krylov complexity, as defined in Eq.(4), for the mass-deformed SYK model as a function of the mass parameter $\kappa$ and the inverse temperature $\beta$. We perform our computations with a system size of $N=26$ and express all dimensionful quantities in units of the coupling $J$, following Ref.García-García et al. (2018). As noted in Balasubramanian et al. (2022), $N=26$ is sufficiently large to ensure convergence of the numerical results. Therefore, we anticipate that finite-size effects on Krylov complexity are minimal for this choice of $N$.
@@
Using the empirical function from Eq.(10) and taking $\kappa_{c}(0)\approx 66$, we estimate that for $\kappa=1$, the critical value of $\beta$ separating chaotic and integrable phases is approximately $\beta_{c}\approx 6.58$. This value lies in between the blue and purple lines in Fig. 3. Our numerical results show that this prediction, derived from the new KCP order parameter, matches the observed disappearance of the linear ramp in SFF$(t)$. This agreement not only verifies that the linear ramp and corresponding dip are indicative of quantum chaotic systems, disappearing in integrable ones, but also independently supports the validity and utility of the order parameter $\Delta C$ as a reliable benchmark for quantum chaotic behavior.
@@ Acknowledgements.
We would like to thank Antonio M. Garcia-Garcia for collaboration at the initial stage of this project and Junggi Yoon for valuable suggestions. MB would also like to thank Dario Rosa for illuminating discussions on complexity and quantum chaos. MB and KBH acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). M.B. acknowledges the sponsorship from the Yangyang Development Fund. HSJ and JFP are supported by the Spanish MINECO ‘Centro de Excelencia Severo Ochoa’ program under grant SEV-2012-0249, the Comunidad de Madrid ‘Atracción de Talento’ program (ATCAM) grant 2020-T1/TIC-20495, the Spanish Research Agency via grants CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033, and ERDF A way of making Europe. KYK was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT &\&& Future Planning (NRF-2021R1A2C1006791) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2024. KYK was also supported by the Creation of the Quantum Information Science R&\&&D Ecosystem (Grant No. 2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). All authors contributed equally to this paper and should be considered as co-first authors.

### v2

@@ Abstract
Modern explorations of quantum chaos have prominently featured both level statistics and OTOCs, bolstered by intriguing links between many-body chaos and quantum gravitational systems Shenker and Stanford (2014, 2015); García-García and Verbaarschot (2016); Cotler et al. (2017a); de Boer et al. (2018); Stanford and Witten (2020). In this context, Krylov complexity Parker et al. (2019); Balasubramanian et al. (2022) has recently emerged as a new valuable tool for characterizing quantum chaos, providing an alternative diagnostic beyond traditional analyses. It has been employed in RMT Balasubramanian et al. (2022, 2023); Tang (2023); Caputa et al. (2024); Bhattacharjee and Nandy (2024) and many other quantum chaotic systems, including quantum billiards Hashimoto et al. (2023); Camargo et al. (2024a); Balasubramanian et al. (2024), spin chains Rabinovici et al. (2022a); Scialchi et al. (2024a); Gill et al. (2024); Bhattacharya et al. (2024); Camargo et al. (2024b); Scialchi et al. (2024b), and various flavors of the Sachdev–Ye–Kitaev (SYK) model Rabinovici et al. (2021); Bhattacharjee et al. (2023a); Hörnedal et al. (2022); Erdmenger et al. (2023a); Chapman et al. (2024). Additionally, Krylov complexity has been discussed in several other contexts including topological and quantum phase transitions Caputa and Liu (2022); Afrasiar et al. (2022); Caputa et al. (2023); Pal et al. (2023), quantum batteries Kim et al. (2022), bosonic systems describing ultra-cold atoms Bhattacharyya et al. (2023), saddle-dominated scrambling Bhattacharjee et al. (2022); Huh et al. (2024), and open quantum systems Bhattacharya et al. (2022); Bhattacharjee et al. (2023b); Mohan (2023); Bhattacharya et al. (2023); Bhattacharjee et al. (2024); Carolan et al. (2024) among others – see Nandy et al. (2024) for a comprehensive review. Two forms of Krylov complexity have been proposed in the literature: the original type, which addresses operator growth Parker et al. (2019), and a newer version, which evaluates the spread of a time-evolving quantum state within a specific subspace of the Hilbert space Balasubramanian et al. (2022). The latter will be the focus of this Letter.
@@
To provide evidence for our proposal, we turn to the SYK model and its variants, which serve as useful toy models. Specifically, we focus on the mass-deformed SYK model Song et al. (2017); Eberlein et al. (2017); García-García et al. (2018), which has been extensively studied for its relevance to quantum chaotic to integrable transitions, and has been thoroughly characterized through both level statistics and properties of the OTOCs García-García et al. (2018); Nosaka et al. (2018); Kim and Cao (2021a); García-García et al. (2021); Lunkin et al. (2020); Nandy et al. (2022); Menzler and Jha (2024).
@@
Given the definitions provided in Eq.(4) and Eq.(6), it is natural to assume that a direct link between Krylov complexity and the SFF might exist. In fact, considering the Krylov complexity of the TFD state, the SFF can be interpreted as the survival probability of the time-evolved TFD state del Campo et al. (2017); Caputa et al. (2024): $\textrm{SFF}(t)=|\psi_{0}(t)|^{2}$. Additionally, for the TFD state with $\beta=0$, which is a maximally-entangled state, the late-time behavior of the Krylov complexity obeys the following identity Cotler et al. (2017a); Rabinovici et al. (2021, 2022b); Erdmenger et al. (2023a)
@@
Results. We calculate the time-dependent Krylov complexity, as defined in Eq.(4), for the mass-deformed SYK model as a function of the mass parameter $\kappa$ and the inverse temperature $\beta$. We perform our computations with a system size of $N=26$ and express all dimensionful quantities in units of the coupling $J$, following Ref.García-García et al. (2018). As noted in Balasubramanian et al. (2022), $N=26$ is sufficiently large to ensure convergence of the numerical results. Therefore, finite-size effects on Krylov complexity are minimal for this choice of $N$. For further details on the numerical methods, we refer the reader to the Supplemental Material.
@@
Using the empirical function from Eq.(10) and taking $\kappa_{c}(0)\approx 66$, we estimate that for $\kappa=1$, the critical value of $\beta$ separating chaotic and integrable phases is approximately $\beta_{c}\approx 6.58$. This value lies in between the blue and purple lines in Fig. 3. Our numerical results show that this prediction, derived from the new KCP order parameter, matches the observed disappearance of the linear ramp in SFF$(t)$. This agreement not only verifies that the presence of a linear ramp and corresponding dip are clear indicators of quantum chaotic systems, which disappear as the system becomes integrable, but it also independently supports the validity and utility of the order parameter $\Delta C$ as a reliable benchmark for assessing quantum chaotic behavior.
@@ Acknowledgements.
We would like to thank Antonio M. Garcia-Garcia for collaboration at the initial stage of this project and Junggi Yoon for valuable suggestions. We would like to thank Pratik Nandy and Debodirna Ghosh for useful comments on a first version of this manuscript. MB would also like to thank Dario Rosa for illuminating discussions on complexity and quantum chaos. MB and KBH acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). M.B. acknowledges the sponsorship from the Yangyang Development Fund. HSJ and JFP are supported by the Spanish MINECO ‘Centro de Excelencia Severo Ochoa’ program under grant SEV-2012-0249, the Comunidad de Madrid ‘Atracción de Talento’ program (ATCAM) grant 2020-T1/TIC-20495, the Spanish Research Agency via grants CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033, and ERDF A way of making Europe. KYK was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT &\&& Future Planning (NRF-2021R1A2C1006791) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2024. KYK was also supported by the Creation of the Quantum Information Science R&\&&D Ecosystem (Grant No. 2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). All authors contributed equally to this paper and should be considered as co-first authors.
Supplemental material
In this Supplementary Material, we offer additional details on the numerical computations discussed in the main text and present further analyses to corroborate our findings.
I Block diagonalization of the SYK Hamiltonian
Block diagonalization of the Hamiltonian matrix is crucial for analyzing spectral statistics of the energy spectrum. By decomposing the Hamiltonian into smaller blocks based on the system’s symmetries, the statistical properties of the energy levels can be studied more efficiently. This approach simplifies computations, improves numerical accuracy, and reduces the effective size of the system.
Regarding the SYK model, its Hamiltonian possesses a conserved charge parity operator $P$ You et al. (2017); Cotler et al. (2017b); Krishnan et al. (2018),
$P=\left\{\begin{array}[]{ll}(-i\chi_{1}\chi_{2})(-i\chi_{3}\chi_{4})\,\cdots\,(-i\chi_{N-1}\chi_{N})&\quad N\in\text{even}\,,\cr(-i\chi_{1}\chi_{2})(-i\chi_{3}\chi_{4})\,\cdots\,(-i\chi_{N}\chi_{\infty})&\quad N\in\text{odd}\,,\end{array}\right.$ (13)
where $\chi_{i}\,(i=1,2,\dots,N)$ are Majorana fermions, and $N$ is the system size. The SYK Hamiltonian $H$ commutes with the parity operator $P$, $[H,P]=0$, allowing the Hamiltonian to be block diagonalized. Using an invertible matrix consisting of the eigenvectors of $P$, the system is split into parity-even and parity-odd sectors, each with a dimension $d=2^{\frac{N}{2}-1}$. Fig. 4 illustrates the typical block diagonalized Hamiltonian of a mass-deformed SYK model at finite $\kappa$.
Figure 4: The block diagonalized mass-deformed SYK model for $N=26$ and $\kappa=1$. The top-left block corresponds to the parity-odd sector, while the bottom-right block corresponds to the parity-even sector.
In our study, we have focused on numerical computations within the parity-even sector for $N=26$, using both Wolfram Mathematica and Matlab to cross-verify our results. We have also confirmed that the parity-odd sector produces similar outcomes for Krylov complexity.
II Lanczos coefficients and quantum chaos
In the main text, we focused on the Krylov complexity derived from solving the Schrödinger equation using the given Lanczos coefficients $\{a_{n},\,b_{n}\}$. Here, we examine the Lanczos coefficients $\{a_{n},\,b_{n}\}$ of the mass-deformed SYK models. As suggested in references Hashimoto et al. (2023); Parlett (1998), the Lanczos coefficients can be obtained through the Lanczos algorithm, which minimizes numerical errors in the orthogonalization process, ensuring consistency with those derived from the Hessenberg form. The algorithm is outlined as follows:
  1. 1.
    Initialize with $b_{0}:=0$, $|K_{0}\rangle:=|\psi(0)\rangle$, $a_{0}:=\langle K_{0}|\mathcal{D}|K_{0}\rangle$, where $\mathcal{D}:=\text{diag}\left(E_{1},\ldots,E_{d}\right)$.
  2. 2.
    For $n\geq 1$: Compute $|A_{n}\rangle=\left(\mathcal{D}-a_{n-1}\right)|K_{n-1}\rangle-b_{n-1}|K_{n-2}\rangle$.
  3. 3.
    Replace $|A_{n}\rangle\rightarrow|A_{n}\rangle-\sum_{m=0}^{n-1}\langle{A_{m}}|{K_{0}}\rangle|A_{m}\rangle$.
  4. 4.
    Set $b_{n}=\langle{A_{n}}|{A_{n}}\rangle^{1/2}$.
  5. 5.
    If $b_{n}=0$, terminate the algorithm; otherwise, set $|{K_{n}}\rangle=b_{n}^{-1}|{A_{n}}\rangle$ and $a_{n}=\langle{K_{n}}|\mathcal{D}|{K_{n}}\rangle$, then return to step 2.
In our study, we utilize the TFD state as the initial state $|\psi(0)\rangle$. To illustrate our numerical results, we present specific computations for the maximally entangled state, i.e., the TFD state with $\beta=0$. Using the Lanczos algorithm, we obtain the Lanczos coefficients for the mass-deformed SYK model, as shown in Fig. 5 below. We observe that $a_{n}$ exhibits oscillatory behavior, while $b_{n}$ initially grows (see inset), reaches a peak, and then diminishes as $n$ approaches the dimension of the Krylov space. These characteristics are consistent with those observed in the original SYK model ($\kappa=0$) Balasubramanian et al. (2022) and in RMT Erdmenger et al. (2023a).
Figure 5: Lanczos coefficients $\{a_{n},\,b_{n}\}$ for the mass-deformed SYK model for various values of the mass-deformation parameter, ranging from $\kappa=0\,(\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{red}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$ to $\kappa=300\,(\color[rgb]{0.58,0.0,0.83}\definecolor[named]{pgfstrokecolor}{rgb}{0.58,0.0,0.83}\textbf{purple}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$.
Two notable observations regarding the effect of $\kappa$ are as follows. First, as $\kappa$ increases, the slope of the initial growth of $b_{n}$ also increases (see the inset), indicating that the value of $b_{1}$ is enhanced by the mass deformation parameter. This enhancement may explain the behavior of the slope of $C(t)$ in Fig. 1 with increasing $\kappa$, as the early-time behavior of the Krylov complexity has been shown to follow $C(t\ll 1)\approx b_{1}^{2}t^{2}$ Huh et al. (2024). The second observation concerns the variance of the Lanczos coefficients, defined as Hashimoto et al. (2023)
$\begin{split}\sigma_{a}^{2}&:=\text{Var}\left(x_{i}^{(a)}\right)\,,\quad x_{i}^{(a)}:=\log\left(\frac{a_{2i-1}}{a_{2i}}\right)\,,\\ \sigma_{b}^{2}&:=\text{Var}\left(x_{i}^{(b)}\right)\,,\quad x_{i}^{(b)}:=\log\left(\frac{b_{2i-1}}{b_{2i}}\right)\,.\end{split}$ (14)
For additional information on the variance of Lanczos coefficients in the context of Krylov operator complexity, see Rabinovici et al. (2022a, b). We observe that the variance increases in the integrable regime (large $\kappa$) compared to the chaotic regime (small $\kappa$). This trend is especially pronounced for $\sigma_{b}^{2}$; see Fig. 6. However, unlike the KCP, we do not observe a distinct feature indicative of a critical phase transition around $\kappa_{c}\approx 66$. This suggests that the KCP may serve as a more effective order parameter than the Lanczos coefficients themselves. Refer to Fig. 7 for the histogram plot of the Lanczos coefficients’ distribution, which supports the conclusions drawn from the variance analysis.
Figure 6: The variance of the Lanczos coefficients $\sigma_{a,b}^{2}$ as a function of $\kappa$.
Figure 7: Histogram of the Lanczos coefficients $a_{n}$ (left panel) and $b_{n}$ (right panel) for values of $\kappa$ ranging from $\kappa=0\,(\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{red}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$ to $\kappa=300\,(\color[rgb]{0.58,0.0,0.83}\definecolor[named]{pgfstrokecolor}{rgb}{0.58,0.0,0.83}\textbf{purple}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$.
III Normalization condition and Ehrenfest theorem
We have confirmed that our numerical results meet essential consistency checks, including the wave function normalization condition Balasubramanian et al. (2022) and the Ehrenfest theorem Erdmenger et al. (2023b).
First, we address the normalization condition for the Krylov wave functions,
$\sum_{n}|\psi_{n}(t)|^{2}=1\,,$ (15)
which ensures the unitarity of time evolution. As shown in Fig. 8, this normalization condition is satisfied within our time window for numerical computations of Krylov complexity, specifically for $t\leq 2^{N/2+1}$. To maintain normalization over a longer time window, including the late-time regime, it is necessary to consider larger values of $n_{\text{max}}$, as discussed in Hashimoto et al. (2023); Camargo et al. (2024a). Our results indicate that the chosen cutoff value of $n_{\text{max}}=d$ is adequate for capturing all key features of Krylov complexity —the initial ramp, peak, decline, and plateau— while satisfying the normalization condition.
Figure 8: The normalization condition of the Krylov wave functions.
Second, an important aspect of Krylov complexity is its adherence to the Ehrenfest theorem Erdmenger et al. (2023a), expressed as:
$\begin{split}\partial^{2}_{t}C(t)=-\left[\left[C(t)\,,\mathcal{L}\right],\mathcal{L}\right]\,,\end{split}$ (16)
where $\mathcal{L}=H\otimes\mathbb{I}$ denotes the Liouvillian and $\mathbb{I}$ represents the identity operator. By applying the Schrödinger equation along with the definition of Krylov complexity, the Ehrenfest theorem (16) can be formulated in terms of the Lanczos coefficients and the Krylov wave functions as follows:
$\begin{split}\partial^{2}_{t}C(t)=2\sum_{n}\left[\left(b_{n+1}^{2}-b_{n}^{2}\right)\psi_{n}(t)\psi_{n}^{*}(t)+\left(a_{n+1}-a_{n}\right)b_{n+1}\psi_{(n+1}(t)\psi_{n)}^{*}(t)\right]\,,\end{split}$ (17)
where $\mathcal{T}_{(a}\tilde{\mathcal{T}}_{b)}:=\frac{1}{2}\left(\mathcal{T}_{a}\tilde{\mathcal{T}}_{b}+\mathcal{T}_{b}\tilde{\mathcal{T}}_{a}\right)$. In Fig. 9, we validate the Ehrenfest theorem for mass-deformed SYK models, demonstrating the relationship between the second time derivative of Krylov complexity and a combination of Lanczos coefficients. It is important to emphasize that this relationship, as expressed in (17), holds universally for any system by construction. This verification strengthens the reliability of our numerical results.
[…]